OpenStudy (anonymous):

A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mathstudent55

OpenStudy (amistre64):

multiply one of them by a useful value, and add it to the other .. a useful value will be such that you produce a negative element of the other one that will cancel out when added

OpenStudy (anonymous):

can you give me an example im still a little confused :/

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

3x+8y=12 ...(1) 2x+2y=3 ...(2) multiply (2) by -4 and add in (1) 3x+8y-8x-8y=12-12 -5x=0 x=0 put in (1) y =?